A 5.0 kg drone is flying perfectly horizontally above the earth’s surface at a height of 40 m and a speed of 15 m/s. It then drops down to a height of 25 m (still horizontally) and reduces its speed to 6 m/s. What is the amount of work that is done on the drone to change its mechanical energy in this way?
1
Expert's answer
2020-05-08T16:28:20-0400

As per the question,

mass of the drone "(m) = 5kg"

Height of the drone "(h_1)=40m"

speed of the drone "(u)=15m\/sec"

final height "(h_2)=25m"

final speed "(v)= 6m\/sec"

So, work done "(w)= mgh_1- mgh_2+\dfrac{m\times u^2}{2}-\dfrac{mv^2}{2}"

"=(5\times 9.8\times 40 - 5\times 9.8\times25 +\dfrac{5\times 15^2}{2}-\dfrac{5\times 6^2}{2})J"


"=(735+472.5 )J"

"=1207.5J"